# Labelling Data with Unknown References

### v2

@@ 8.1.1 Ablation: Lying Evaluators and Lying Generators
Table 3: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.6$. From top to bottom: Lie 1: $\sigma$ unknown. Lie 2: $f$ unknown. Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step. In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$. Note the comparatively high success rates in the IP case of Lies 1 and 2. We discuss these findings in Section 9. Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
8.1.2 Ablation: LLM Generation Strategies
Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol. Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [30]. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 8.1, and actually generating a datapoint from the rubric. The results are in Table 4. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation led to bad performance in both LLMs. We note that this is expected: in general, classifying an element in a language is much easier than generating an element [17, 16].
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
Table 4: Results of our ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline. We report the highest-performing run at $\phi=0.9$. While both models were able to broadly adjust to their original accuracy, their success rates are markedly different when the models are generating a datapoint versus picking one.
9 Discussion
9.1 Theory
Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. However, this result has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue this interpretation also makes sense: otherwise, it would be possible to create information out of thin air (that is, without any prior knowledge of $f$). Moreover, the primary purpose of the No-Data Algorithm is to establish trustworthiness on the evaluator, not label the data.
One important limitation of the theory is that we assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings. The first assumption can be addressed by ‘binarising’ a set $Y$ with arity $>2$ by separating it into two classes. The second, however, must be understood more carefully: if this were a natural-language problem, it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings. Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual.
9.2 Experiments
The results adjusted to the theory’s predictions. In particular, in both the known/unknown comparison and the lying ablation study, the LLM and DT evaluators consistently had low success rates for the OOP case, thus indicating that the No-Data Algorithm is robust to evaluator strategies attempting deception. Three things we noted throughout our experimentation were the high number of flips in the DT OOP case; the (relatively) high success rate when the evaluator lied about knowing $\sigma$; and the low success rate of the evaluator with approximate knowledge of $f$.
The first illustrates our interpretation of Theorem 1: since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie. Since the flip probability is set to $\phi=0.5$, half the labels got flipped and no information was gained from the labelling. However, we learnt that the evaluator cannot be trusted–the main point of the No-Data Algorithm.
In the case of the evaluator lying about $\sigma$, the explanation is simple. This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2). However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability. In our experiments, the internal structure depends on both clauses in $c_{2}$.
Finally, the (relatively) low success rate of the evaluator with approximate knowledge of $f$ illustrates one critical characteristic of the No-Data Algorithm: the success rate is a very sensitive metric, highly dependent on both the evaluator’s labelling and generation reliabilities.
10 Conclusion
We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist. It works by performing checks that we showed to be sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels. Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios. Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted.
One natural implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated. We argued that this scenario is unsolvable: there is no prior information on the datapoint, and hence it will be effectively creating information out of thin air. This is why the algorithm’s focus is less on the labelling returned, and more on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments). Alternatively, high accuracy but low success rate indicates deception (r. the OOP case). The addition of the flip probability $\phi$ and the success count allows for some–limited–heuristics by which the No-Data Algorithm can produce a labelling when the evaluator’s output is rejected. Nonetheless, since the primary purpose of this algorithm is to establish trustworthiness, and not to label the set, it is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
As mentioned, some of our work’s limitations are solvable (e.g., label sets of arity $>$ 2), and others are natural (the interpretation of Theorem 1). However, other limitations are more practical in nature and should be the subject of further work. Namely, it is well-known that LLMs are sensitive to prompting. Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model. This nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
11 Limitations
Most of the limitations in the theory are covered in Section 9, and relevant extensions to more common setups are provided in Appendix C. Even when our focus was not experimentation, it could be argued that these were limited in size and scope. We agree–time constraints made us prioritise rigour over experimental results. We have open-sourced the code and data for our work, but also plan to extend it ourselves with other datasets and open models in the near future.
12 Ethics
We argue that the nature of our work precludes any misuse.
Appendix A Detailed Proofs
A.1 Correctness of the EV protocol
A.1.1 Part 1 (Correctness)
We prove the following lemma:
Lemma 2.
Proof.
(Sufficient) It follows directly from the earlier statement. Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $x\not\equiv x^{\prime}$, $x\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$. This would imply that (1) either the string , a contradiction.
This concludes the proof. ∎
A.1.2 Part 2 (Robustness)
We prove the following lemma:
Lemma 3.
Proof.
In here we have two scenarios: either the evaluator knows $f$, or it is lying or cheating.
A.2 Lie 1: Generate a ‘good’ $x^{\prime}$
A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
Therefore, from Lemma 2 we know that both challenges together are sufficient to ascertain the truthfulness of a label, and from Lemma 3 that the ability of the verifier to do this is bounded. Putting both lemmas together we obtain the proof for Lemma 1.
A.4 Correctness Bounds of the No-Data Algorithm
For this, it is much easier to work with the errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
This yields:
[…]
$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$. (9)
Appendix B Characterisation Bounds
Lemma 4.
Suppose the evaluator and verifier run in polynomial time for any input $x\in X$. Then, if the No-Data Algorithm is ran over a dataset $D$; and with $r$ rounds of EV protocol, the runtime is then $\operatorname{O}\!\left(r\lvert D\rvert\right)$.
Proof.
Straightforward. ∎
Appendix C Extensions
C.1 Distance-based Constraints
C.2 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm $k-1$ times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of $k-1$.
Appendix D Detailed Experimental Methods
D.1 Detailed Setup
@@
For the LLM we used o3-mini (version 2025-01-31), and called it via the Azure OpenAI API. We set the maximum tokens to 50,000 and left all other parameters as default. For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
D.2 Prompts
@@
The prompts we used for our LLM stack are in Prompt 1 (evaluation) and Prompt 2 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string.
@@
<NL RUBRIC GOES HERE>
@@
Table 1: System prompt for evaluation. Anchors (‘|reasons|’, ‘|label|’) are used for parsing. The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
@@
<NL RUBRIC GOES HERE>
@@
Table 2: System prompt for generation. This prompt does not utilise exemplars. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric and selected datapoints from the IP training dataset, ensuring that at least one of the entries matched the criterion. Given that the model did not follow the output format exactly, some trial-and-error was needed to ensure a comparatively high parse success rate.
@@
Table 3: System prompt for generation used in the ablation study from Section 8.1.2. Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.

### v3

@@ D.2 Ablation: LLM Generation Strategies
Table 5: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.4$. From top to bottom: Lie 1: $\sigma$ unknown. Lie 2: $f$ unknown. Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step. In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$. Note the comparatively high success rates in the IP case of Lies 1 and 2. We discuss these findings in Section 7. Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
D.3 Ablation: Linguistic Competency
In this section we present an expanded version of the results from Section 6.2, including the performance of Qwen 2.5B VL 7B on the same corpus, with GPT-4.1 as a verifier. This model has no indications of being proficient in West Frisian, although the calibration of GPT-4.1 in IP indicates that the verifier is capable of acting as a reliable component. Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate was still low. In the OOP case (where it was set up to lie), the success rate was lowest. As before, this number remained steady throughout the experimentation. The full results are in Table 6.
IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
Known
Predicting $y$ 76.3 / 80.3 51.5 / 67.2 –/– –/–
Avg. per $c\in\bar{C}$ 89.5 / 94.0 45.3 / 46.5 –/– –/–
Unknown
GPT-4.1
$\phi=0.7$ 72.4 / 76.6 49.3 / 57.8 87.8 / 9. 3 1.4 / 70.3
$\phi=0.3$ 74.4 / 78.1 49.1 / 39.1 86.8 / 4.9 1.2 / 33.6
$\phi=0.1$ 76.9 / 79.8 50.7 / 16.4 86.2 / 1.8 1.2 / 9.1
Qwen 2.5B
$\phi=0.7$ 62.5 / 63.9 51.7 / 59.5 34.4 / 46.0 1.8 / 70.3
$\phi=0.3$ 55.9 / 64.8 51.5 / 43.2 34.2 / 21.4 2.1 / 34.4
$\phi=0.1$ 55.5 / 67.7 49.7 / 19.8 35.0 / 6.2 2.1 / 11.5
Table 6: Performance of the LLM on the natural-language version of the No-Data Algorithm with our ablation study including a model not proficient in the tasks. The success rate for this model was low in both the IP case (the model not lying; just wrong) and the OOP case (the model lying and wrong), although much lower in the latter.
Appendix E Rubrics
We present the full rubrics used in our experiments. Rubrics 2 and 3 have the specifications for the natural-language experiments. Both rubrics were passed verbatim to the LLM and used during annotation.
Criterion Definition
c1) The response must be in West Frisian.
c2a) The response must be culturally (e.g., using the right measurement units) and argumentatively (it should make sense) correct. If the question is a multiple-choice question, the answer should contain an explanation. If it requests code, it should also contain an explanation that is clear. Grammar or accuracy of the response are not measured here.
c2b) The response must be correct. If it is code, it should not have syntax errors.
c3) The response must be grammatically correct: coherent, good spelling, etc. Code syntax is not measured here.
c4) The response must not be cut off.
c5) The model must follow the instructions from the user (the prompt) exactly and completely, even if its answer is wrong. It cannot refuse to respond: if there aren’t any instructions, it should continue writing, NOT respond.
Table 2: Rubric for the natural-language evaluation (IF) experiment. This rubric is meant to evaluate the quality of the output of an LLM. The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
Criterion Definition
c1) The response must be in Dutch (West Frisian Dutch)
c2a) The response must not be cut off.
c2b) The response must make as many references as possible to Dutch culture.
c3) The response must continue writing if there is no prompt.
c4) (A) The response must not be cut off, or (B) The response must make as many references as possible to Dutch culture. At least one of them must be correct. If they are both zero or one, the response is zero.
c5) The response must provide a summary or conclusion. It must be explicitly marked as ‘summary’ or ‘conclusion’.
Table 3: Rubric for the natural-language evaluation (OOF). This rubric was designed to be somewhat orthogonal to the IF version. That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IF, c2a here is c4 in IF). Some criteria are completely different, such as c1 and c5.
Appendix F Detailed Experimental Methods
@@
For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default. GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero. DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace. For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters.
We carried out the second experiment (natural language) with GPT-4.1 (version longco-2025-04-14). The evaluator (labelling, criteria) and verifier were set at 256 maximum tokens, while the generator (i.e., the evaluator generating similar datapoints) was set at 5,000 tokens. All models were called with temperature set to zero.
LLMs were sometimes prone to return unparseable outputs. To mitigate this, in case of failure, we re-queried the models up to five times.
All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
F.1 Prompts
@@
The prompts we used for our LLM evaluation on synthetic experiments are in Prompt 1 (evaluation) and Prompt 2 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string. For the natural language experiments the prompts are in Prompt 4 (evaluator/verifier) and Prompt 5 (generator; part of the evaluator). These prompts required no exemplars. When parsing errors occurred, we retried up to five times, and otherwise returned a zero-labelled value.
@@
<NATURAL-LANGUAGE RUBRIC GOES HERE>
@@
Table 1: System prompt for the evaluator (synthetic experiment). Anchors (‘|reasons|’, ‘|label|’) are used for parsing. The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
@@
<NATURAL-LANGUAGE RUBRIC GOES HERE>
@@
Table 2: System prompt for the generator (synthetic experiment). This prompt does not utilise exemplars. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric and selected datapoints from the IP training dataset, ensuring that at least one of the entries matched the criterion. Given that the model did not follow the output format exactly, some trial-and-error was needed to ensure a comparatively high parse success rate.
@@
Table 3: System prompt for generation used in the ablation study from Appendix D.2. Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
You are an LLM evaluator. You will be given a prompt and an response in West Frisian, meant for West Frisian readers.
Your job will be to verify if the response follows certain criteria and give a final binary score.
Check the output against the criteria below. If it fulfils the criteria, it should be a 1. Otherwise, 0.
If any of the criteria score a zero, the response must be zero.
# Criteria:
<CRITERIA GO HERE>
# Output format:
<OUTPUT FORMAT GOES HERE>
Table 4: System prompt for the natural-language evaluator and verifier. The output format requested is always in JSON. Depending on the setup, the output format and criteria will vary. For example, if it is meant to evaluate a single criterion, say, c2b, the definition will include only the definition of said criterion and the schema will specify solely ‘c2b’ as a key. Remark that the verifier will always expect the criteria from the IF rubric.
You are a paraphraser evaluating a prompt and an output for an LLM.
You will be given a datapoint (prompt/output), a label, and a list of reasons why that datapoint’s output has that label.
Your job will be to return a SIMILAR prompt and output, such that the OUTPUT (1) it matches the list of reasons, and (2) matches the label.
The output must match the values in the list of reasons.
Here’s the rubric used for these reasons:
<RUBRIC GOES HERE>
Your response must be in JSON using the following schema:
 "Prompt": the new, paraphrased user prompt.
 "Output": the new, paraphrased output fulfiling the criteria.
Only use the keys "Prompt" and "Output"
Table 5: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IF or OOF), as labelled by the evaluator itself.